# Multivariate Probabilistic Time Series Forecasting with Correlated Errors

### v3

@@ Abstract
Accurately modeling the correlation structure of errors is essential for reliable uncertainty quantification in probabilistic time series forecasting. Recent deep learning models for multivariate time series have developed efficient parameterizations for time-varying contemporaneous covariance, but they often assume temporal independence of errors for simplicity. However, real-world data frequently exhibit significant error autocorrelation and cross-lag correlation due to factors such as missing covariates. In this paper, we present a plug-and-play method that learns the covariance structure of errors over multiple steps for autoregressive models with Gaussian-distributed errors. To achieve scalable inference and computational efficiency, we model the contemporaneous covariance using a low-rank-plus-diagonal parameterization and characterize cross-covariance through a group of independent latent temporal processes. The learned covariance matrix can be used to calibrate predictions based on observed residuals. We evaluate our method on probabilistic models built on RNN and Transformer architectures, and the results confirm the effectiveness of our approach in enhancing predictive accuracy and uncertainty quantification without significantly increasing the parameter size.
@@ 1 Introduction
Uncertainty quantification is vital in time series forecasting, especially for applications that need more detailed insights than point forecasts. Probabilistic time series forecasting with deep learning (DL) has gained attention for its ability to capture complex/nonlinear dependencies and provide the probability distribution of target variables [1]. In multivariate time series, autoregressive models are popular for probabilistic forecasting [2, 3, 4], modeling the joint one-step-ahead predictive distribution and making multistep-ahead predictions in a rolling manner. To achieve scalable learning, these models often assume that errors are independent over time. Typically, time series variables follow a Gaussian distribution $\mathbf{z}_{t}=m(\mathbf{h}_{t})+\boldsymbol{\eta}_{t}$, where $m(\cdot)$ is the mean function and $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$ is a stochastic error process with contemporaneous covariance matrix $\boldsymbol{\Sigma}_{t}$. The assumption of time-independence implies $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$, $\forall s\neq t$. This holds when the model can incorporate all correlations between successive time steps via hidden states determined by previous values. However, real-world data often violate this assumption, with residuals showing substantial cross-correlation due to omission of important covariates and model misspecification.
@@
Modeling error autocorrelation (or cross-correlation) is a key research topic in statistical time series models. A common approach is to assume that the error series follows a dependent temporal process, such as an autoregressive integrated moving average (ARIMA) model [5]. DL models face similar challenges. Previous studies have tried to incorporate temporally-correlated errors into the training process by modifying the loss function [6, 7]. However, these methods, based on deterministic output, are not easily applicable to probabilistic forecasting models, especially in a multivariate setting. A notable innovation is the batch training method introduced by Zheng et al. [8], which trains a univariate probabilistic forecasting model using generalized least squares (GLS) loss over batched errors. This approach parameterizes a dynamic covariance matrix to characterize error autocorrelation, which is then used to calibrate the predictive distribution of time series variables. This approach consistently improves probabilistic forecasting performance compared to naive training (i.e., without considering autocorrelated errors); however, it is only applicable to univariate models, such as DeepAR [9].
@@
In this paper, we introduce an efficient method for learning error cross-correlation in multivariate probabilistic forecasting models. We focus on deep learning models that are autoregressive with Gaussian-distributed errors. Modeling cross-correlation in multivariate models is challenging due to increased dimensionality, as the covariance matrix scales with the number of time series $N$. To address this computational challenge, we propose characterizing error cross-correlation with a group of independent latent temporal processes using a low-rank parameterization of the covariance matrix. This approach prevents the computational cost from increasing with the number of time series. Our method results in a general-purpose approach to multivariate probabilistic forecasting models, offering significantly improved predictive accuracy.
@@
    We introduce a plug-and-play method for training autoregressive multivariate probabilistic forecasting models with a re-designed GLS loss. (§4)
@@
    We propose an efficient parameterization of the covariance matrix of errors over multiple steps, allowing for efficient computation of its inverse and determinant using the matrix inversion and determinant lemmas. (§4.1)
@@
    The learned covariance matrix can be used to fine-tune the predictive distribution conditional on observed residuals. (§4.2)
@@
    We demonstrate that the proposed method effectively captures error cross-correlation and improves prediction quality. Importantly, these improvements are realized via a statistical formulation without significantly increasing the size of model parameters. (§5)
@@ 2 Probabilistic Time Series Forecasting
Denote $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$ as the vector containing the time series variables at time step $t$, where $N$ is the number of time series. Probabilistic time series forecasting can be formulated as estimating the joint conditional distribution $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$ based on the observed history $\{\mathbf{z}_{t}\}_{t=1}^{T}$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$ and $\mathbf{x}_{t}$ are known time-dependent covariates (e.g., time of day and day of week) for all future time steps. In essence, the problem can be described as predicting the time series values for $Q$ future time steps, using all available covariates and $P$ steps of historical time series values:
@@
which becomes an autoregressive model that can be used for either one-step-ahead ($Q=1$) or multistep-ahead forecasting in a rolling manner. When performing multistep-ahead forecasting, samples will be drawn in the prediction range ($t\geq T+1$) and fed back for the next time step until the end of the desired prediction range. In neural networks, the conditioning information is commonly encoded into a state vector $\mathbf{h}_{t}$. Hence, Eq. (1) can be expressed more concisely:
@@
Existing autoregressive models usually assume that the error at each time step is independent and thus $\mathbf{z}_{t}$ follows a multivariate Gaussian distribution:
@@
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ map $\mathbf{h}_{t}$ to the mean and covariance parameters of a multivariate Gaussian distribution. This formulation is equivalent to decomposing $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$ with $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. The temporally independent error assumption corresponds to $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$ for any time points $s$ and $t$ such that $s\neq t$. Fig. 1 provides an empirical example of the contemporaneous covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t}),\Delta=1,2,3$. The results are calculated based on the prediction residuals of GPVar [2] on the $\mathtt{m4\_hourly}$ dataset. While multivariate models primarily focus on modeling the contemporaneous covariance, it is evident that the residuals are not independent across time, as $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})\neq\mathbf{0}$. The non-zero cross-covariance suggests that there still exists information left in the residuals, and we can leverage it to enhance the prediction.
@@ 3.1 Probabilistic Time Series Forecasting
Probabilistic forecasting aims to model the probability distribution for target variables, unlike deterministic forecasting which produces only point estimates. There are two main approaches: parametric probability density functions (PDFs) and quantile functions [1]. For instance, MQ-RNN [10] generates quantile forecasts using a sequence-to-sequence (Seq2Seq) RNN architecture. In contrast, PDF-based approaches assume a specific distribution (e.g., Gaussian, Poisson) and use neural networks to generate the distribution parameters. DeepAR [9], for instance, uses an RNN to model hidden state transitions, while its multivariate version, GPVar [2], employs a Gaussian copula to transform observations into Gaussian variables, assuming a joint multivariate Gaussian distribution.
Neural networks can also generate probabilistic model parameters. The deep state space model (SSM) [11] uses an RNN to generate SSM parameters. The normalizing Kalman filter (NKF) [12] combines normalizing flows (NFs) with the linear Gaussian state space model (LGM) to model nonlinear dynamics and evaluate the PDF of observations. NKF uses RNNs to produce LGM parameters at each time step, then transforms the LGM output into observations using NFs. Wang et al. [13] proposed the deep factor model, which includes a deterministic global component parameterized by an RNN and a random component from any classical probabilistic model (e.g., Gaussian white noise) to represent random effects. Some methods improve expressive conditioning for probabilistic forecasting by using Transformers instead of RNNs to model latent state dynamics, thus breaking the Markovian assumption in RNNs [14]. Other approaches adopt more flexible distribution forms, including normalizing flows [3], diffusion models [4], and Copulas [15, 16]. We refer readers to Benidis et al. [1] for a recent and comprehensive review.
@@ 3.2 Modeling Correlated Errors
Error correlation in time series is a well-studied topic in the domains of econometrics and statistics [17, 5, 18]. In multivariate time series, correlation structure is characterized by contemporaneous covariance $\operatorname{Var}(\boldsymbol{\eta}_{t})=\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})$. Cross-covariance includes both the autocovariance of errors $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{i,t})$ and the cross-lag covariance $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{j,t})$ between all pairs of components in the multivariate series. Contemporaneous covariance captures the correlation among individual time series at a specific point in time. In the univariate setting, DeepAR [9] achieves probabilistic forecasting by modeling the contemporaneous covariance, assuming that errors are independent over time. To address autocorrelation, Sun et al. [6] re-parameterized the input and output of neural networks to model first-order error autocorrelation, effectively capturing serially correlated errors using an AR$(1)$ process. This method enhances the performance of one-step-ahead neural forecasting models, allowing joint optimization of base and error regressors, but is limited to deterministic models. In spatial modeling, Saha et al. [19] introduced the RF-GLS model, which uses random forests to estimate nonlinear covariate effects and Gaussian processes (GP) to model spatial random effects. The RF-GLS model assumes that the error process follows an AR$(p)$ process to accommodate autocorrelated errors. Zheng et al. [8] proposed training a probabilistic forecasting model with a GLS loss that explicitly models the time-varying autocorrelation of batched error terms, extending DeepAR to incorporate autocorrelated errors.
@@
In the multivariate setting, most existing work focuses on modeling contemporaneous covariance, assuming that $\boldsymbol{\eta}_{t}$ is independently distributed, which implies $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})=\boldsymbol{0}$. For example, GPVar [2] generalizes DeepAR [9] to account for correlations between time series by viewing the distribution of time series variables as a Gaussian process. In Seq2Seq models, correlations can span across series and forecasting steps. Since predictions for future time steps are generated simultaneously, we refer to these correlations as contemporaneous correlations within the scope of this study. Choi et al. [20] introduced a dynamic mixture of matrix Gaussian distributions to capture contemporaneous covariance of errors in Seq2Seq models. An exception that explicitly models error cross-correlation is [7], where the authors assume that the matrix-variate error term of a multivariate Seq2Seq model follows a matrix autoregressive (AR) process with seasonal lags. However, applying this technique to probabilistic forecasting models is not straightforward.
@@
To the best of our knowledge, our work represents the first attempt to model cross-covariance in multivariate probabilistic time series forecasting. The closest related studies are by Zheng et al. [8] and Zheng et al. [7]. Zheng et al. [8] applies GLS loss in the temporal domain to model autocorrelated errors, but their approach is tailored for univariate time series. Zheng et al. [7] models cross-covariance in multivariate forecasting models, but their method is limited to deterministic models and requires predefined seasonal lags in the error autoregressive process. Our work extends [8] to the multivariate setting, enabling the modeling of the correlation structure of multivariate errors across multiple steps. In addition, we distinguish our approach from methods that directly model the distribution of time series variables, such as Copulas [15, 16], where no decomposition of the error term is provided.

### v4

@@ Abstract
Accurately modeling the correlation structure of errors is critical for reliable uncertainty quantification in probabilistic time series forecasting. While recent deep learning models for multivariate time series have developed efficient parameterizations for time-varying contemporaneous covariance, but they often assume temporal independence of errors for simplicity. However, real-world data often exhibit significant error autocorrelation and cross-lag correlation due to factors such as missing covariates. In this paper, we introduce a plug-and-play method that learns the covariance structure of errors over multiple steps for autoregressive models with Gaussian-distributed errors. To ensure scalable inference and computational efficiency, we model the contemporaneous covariance using a low-rank-plus-diagonal parameterization and capture cross-covariance through a group of independent latent temporal processes. The learned covariance matrix is then used to calibrate predictions based on observed residuals. We evaluate our method on probabilistic models built on RNNs and Transformer architectures, and the results confirm the effectiveness of our approach in improving predictive accuracy and uncertainty quantification without significantly increasing the parameter size.
@@ 1 Introduction
Uncertainty quantification is crucial in time series forecasting, especially for applications that need more detailed insights than point forecasts. Probabilistic time series forecasting with deep learning (DL) has attracted attention for its ability to capture complex, nonlinear dependencies and provide the probability distribution of target variables [1, 2]. In multivariate time series, autoregressive models are widely used for probabilistic forecasting [3, 4, 5], modeling the joint one-step-ahead predictive distribution and generating multistep-ahead predictions in a rolling manner. To enable scalable learning, these models often assume that errors are independent over time. Typically, time series variables follow a Gaussian distribution $\mathbf{z}_{t}=m(\mathbf{h}_{t})+\boldsymbol{\eta}_{t}$, where $m(\cdot)$ is the mean function and $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$ is a stochastic error process with contemporaneous covariance matrix $\boldsymbol{\Sigma}_{t}$. The assumption of time-independence implies $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$, $\forall s\neq t$. This holds when the model can account for all correlations between successive time steps through hidden states determined by previous values. However, real-world data often violate this assumption, as residuals exhibit substantial cross-correlation due to omission of important covariates and model misspecification.
@@
Modeling error autocorrelation (or cross-correlation) is a key area of research in statistical time series models. A common approach is to assume that the error series follows a dependent temporal process, such as an autoregressive integrated moving average (ARIMA) model [6]. Deep learning models face similar challenges. Previous studies have attempted to incorporate temporally-correlated errors into the training process by modifying the loss function [7, 8]. However, these methods, based on deterministic output, are not easily applicable to probabilistic forecasting models, particularly in a multivariate setting. A notable innovation is the batch training method introduced by Zheng et al. [9], which trains a univariate probabilistic forecasting model using generalized least squares (GLS) loss over batched errors. This approach parameterizes a dynamic covariance matrix to capture error autocorrelation, which is then used to calibrate the predictive distribution of time series variables. While this method consistently improves probabilistic forecasting performance compared to naive training (i.e., without considering autocorrelated errors); however, it is only applicable to univariate models, such as DeepAR [10].
@@
In this paper, we introduce an efficient method for learning error cross-correlation in multivariate probabilistic forecasting models. Our focus is on deep learning models that are autoregressive with Gaussian-distributed errors. Modeling cross-correlation in multivariate models presents challenges due to increased dimensionality, as the covariance matrix scales with the number of time series $N$. To address this computational challenge, we propose characterizing error cross-correlation through a set of independent latent temporal processes using a low-rank parameterization of the covariance matrix. This approach prevents the computational cost from growing with the number of time series. Our method offers a general-purpose approach to multivariate probabilistic forecasting models, offering significantly improved predictive accuracy.
@@
    We introduce a plug-and-play method for training autoregressive multivariate probabilistic forecasting models using a redesigned GLS loss. (§4)
@@
    We propose an efficient parameterization of the error covariance matrix across multiple steps, enabling efficient computation of its inverse and determinant through matrix inversion and determinant lemmas. (§4.1)
@@
    The learned covariance matrix is used to fine-tune the predictive distribution based on observed residuals. (§4.2)
@@
    We demonstrate that the proposed method effectively captures error cross-correlation and improves prediction quality. Notably, these improvements are achieved through a statistical formulation without significantly increasing the size of model parameters. (§5)
@@ 2 Probabilistic Time Series Forecasting
Denote $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$ as the vector of time series variables at time step $t$, where $N$ is the number of time series. Probabilistic time series forecasting can be formulated as estimating the joint conditional distribution $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$ given the observed history $\{\mathbf{z}_{t}\}_{t=1}^{T}$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$ and $\mathbf{x}_{t}$ are known time-dependent covariates (e.g., time of day, day of week) for all future time steps. In essence, the problem involves predicting the time series values for $Q$ future time steps using all available covariates and $P$ steps of historical time series data:
@@
which becomes an autoregressive model that can be used for either one-step-ahead ($Q=1$) or multistep-ahead forecasting in a rolling manner. When performing multistep-ahead forecasting, samples are drawn in the prediction range ($t\geq T+1$) and fed back for the next time step until the end of the desired prediction range. In neural networks, the conditioning information is commonly encoded into a state vector $\mathbf{h}_{t}$. Hence, Eq. (1) can be expressed more concisely:
@@
Existing autoregressive models typically assume that the error at each time step is independent, meaning that $\mathbf{z}_{t}$ follows a multivariate Gaussian distribution:
@@
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ map $\mathbf{h}_{t}$ to the mean and covariance parameters of a multivariate Gaussian distribution. This formulation can be decomposed as $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$ with $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. The temporally independent error assumption corresponds to $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$ for any time points $s$ and $t$ where $s\neq t$. Fig. 1 provides an empirical example of the contemporaneous covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t}),\Delta=1,2,3$. The results are calculated based on the prediction residuals of GPVar [3] on the $\mathtt{m4\_hourly}$ dataset. While multivariate models primarily focus on contemporaneous covariance, the residuals clearly exhibit temporal dependence, as $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})\neq\mathbf{0}$. This non-zero cross-covariance suggests that residuals still contain valuable information, which can be leveraged to improve predictions.
@@ 3.1 Probabilistic Time Series Forecasting
Probabilistic forecasting aims to model the probability distribution of target variables, unlike deterministic forecasting, which produces only point estimates. There are two main approaches: parametric probability density functions (PDFs) and quantile functions [2]. For example, MQ-RNN [11] generates quantile forecasts using a sequence-to-sequence (Seq2Seq) RNN architecture. In contrast, PDF-based approaches assume a specific distribution (e.g., Gaussian, Poisson) and use neural networks to generate the distribution parameters. DeepAR [10], for instance, uses an RNN to model hidden state transitions, while its multivariate version, GPVar [3], employs a Gaussian copula to transform observations into Gaussian variables, assuming a joint multivariate Gaussian distribution.
Neural networks can also generate probabilistic model parameters. The deep state space model (SSM) [12] uses an RNN to generate SSM parameters. The normalizing Kalman filter (NKF) [13] combines normalizing flows (NFs) with the linear Gaussian state space model (LGM) to model nonlinear dynamics and evaluate the PDF of observations. NKF uses RNNs to produce LGM parameters at each time step, then transforms the LGM output into observations using NFs. Wang et al. [14] proposed the deep factor model, which includes a deterministic global component parameterized by an RNN and a random component from any classical probabilistic model (e.g., Gaussian white noise) to represent random effects. Some methods improve expressive conditioning for probabilistic forecasting by using Transformer instead of RNNs to model latent state dynamics, thus breaking the Markovian assumption in RNNs [15]. Other approaches adopt more flexible distribution forms, including normalizing flows [4], diffusion models [5], and copulas [16, 17]. For a recent and comprehensive review, we refer readers to Benidis et al. [2].
@@ 3.2 Modeling Correlated Errors
Error correlation in time series has been extensively studied in econometrics and statistics [18, 6, 19]. In multivariate time series, correlation structure is characterized by contemporaneous covariance $\operatorname{Var}(\boldsymbol{\eta}_{t})=\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})$. Cross-covariance includes both the autocovariance of errors $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{i,t})$ and the cross-lag covariance $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{j,t})$ between pairs of components in the multivariate series. Contemporaneous covariance captures the correlation among individual time series at a specific point in time. In the univariate setting, DeepAR [10] achieves probabilistic forecasting by modeling the contemporaneous covariance, assuming that errors are independent over time. To address autocorrelation, Sun et al. [7] re-parameterized the input and output of neural networks to model first-order error autocorrelation, effectively capturing serially correlated errors using an AR$(1)$ process. This method improves the performance of one-step-ahead neural forecasting models, allowing joint optimization of base and error regressors, but is limited to deterministic models. In spatial modeling, Saha et al. [20] introduced the RF-GLS model, which uses random forests to estimate nonlinear covariate effects and Gaussian processes (GP) to model spatial random effects. The RF-GLS model assumes that the error process follows an AR$(p)$ process to accommodate autocorrelated errors. Zheng et al. [9] proposed training a probabilistic forecasting model with a GLS loss that explicitly models the time-varying autocorrelation of batched error terms, extending DeepAR to incorporate autocorrelated errors.
@@
In the multivariate setting, most existing work focuses on modeling contemporaneous covariance, assuming that $\boldsymbol{\eta}_{t}$ is independently distributed, which implies $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})=\boldsymbol{0}$. For example, GPVar [3] generalizes DeepAR [10] to account for correlations between time series by viewing the distribution of time series variables as a Gaussian process. In Seq2Seq models, correlations can span across series and forecasting steps, as predictions for future time steps are generated simultaneously. Since predictions for future time steps are generated simultaneously, we refer to these correlations as contemporaneous correlations within the scope of this study. Choi et al. [21] introduced a dynamic mixture of matrix Gaussian distributions to capture contemporaneous covariance of errors in Seq2Seq models. One exception that explicitly models error cross-correlation is [8], where the authors assume that the matrix-variate error term of a multivariate Seq2Seq model follows a matrix autoregressive (AR) process with seasonal lags. However, applying this technique to probabilistic forecasting models is not straightforward.
@@
To the best of our knowledge, our work is the first to model cross-covariance in multivariate probabilistic time series forecasting. The closest related studies are by Zheng et al. [9] and Zheng et al. [8]. Zheng et al. [9] applies GLS loss in the temporal domain to model autocorrelated errors, but their approach is tailored for univariate time series. Zheng et al. [8] models cross-covariance in multivariate forecasting models, but their method is limited to deterministic models and requires predefined seasonal lags in the error autoregressive process. Our work extends [9] to the multivariate setting, enabling the modeling of the correlation structure of multivariate errors across multiple steps. In addition, we distinguish our approach from methods that directly model the distribution of time series variables, such as Copulas [16, 17], where no decomposition of the error term is provided.